OpenStudy (anonymous):

Which statement is an example of the Reflexive Property of Equality? A. If p = q, then p – s = q – s. B. x + a = x + b C. x = x D. If p = q, then p + s = q + s.

OpenStudy (anonymous):

It would appear to be D, if p = q, then p + s = q + s.

OpenStudy (anonymous):

I need to fix A.

OpenStudy (anonymous):

Reflexive Property states that for every real number x, x = x.

OpenStudy (anonymous):

An example of reflexive property is CD=CD so what answer looks like that example

OpenStudy (anonymous):

The Symmetric Property states that for all real numbers x and y, if x = y, then y = x.

OpenStudy (anonymous):

Oh my mistake, C would be the closest answer.

OpenStudy (anonymous):

A = If p=q, then p-s = q-s

OpenStudy (anonymous):

Yes, so to make A an example of the reflexive property of equality, you might write If p = q, then q = p.

OpenStudy (anonymous):

but yes I see, x=x just like what you said as ex: CD=CD

OpenStudy (anonymous):

Thank you all.

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

I have one more question. Could someone give me a easy example of Law of Detachment?

OpenStudy (anonymous):

easest one I could find